### v2

# FB-Bench: A Fine-Grained Multi-Task Benchmark for Evaluating LLMs’ Responsiveness to Human Feedback

Abstract
Human feedback is crucial in the interactions between humans and Large Language Models (LLMs). However, existing research primarily focuses on benchmarking LLMs in single-turn dialogues. Even in benchmarks designed for multi-turn dialogues, the user inputs are often independent, neglecting the nuanced and complex nature of human feedback within real-world usage scenarios. To fill this research gap, we introduce FB-Bench, a fine-grained, multi-task benchmark designed to evaluate LLMs’ responsiveness to human feedback under real-world usage scenarios in Chinese. Drawing from the two main interaction scenarios, FB-Bench comprises 591 meticulously curated samples, encompassing eight task types, five deficiency types of response, and nine feedback types. We extensively evaluate a broad array of popular LLMs, revealing significant variations in their performance across different interaction scenarios. Further analysis indicates that task, human feedback, and deficiencies of previous responses can also significantly impact LLMs’ responsiveness. Our findings underscore both the strengths and limitations of current models, providing valuable insights and directions for future research. Code and datasets are available at https://github.com/PKU-Baichuan-MLSystemLab/FB-Bench.
FB-Bench: A Fine-Grained Multi-Task Benchmark for Evaluating LLMs’ Responsiveness to Human Feedback
Youquan Li1, Miao Zheng2, Fan Yang2, Guosheng Dong2, Bin Cui1, Weipeng Chen2, Zenan Zhou2thanks: Corresponding author., Wentao Zhang [^1] footnotemark: 1 1 Peking University 2 Baichuan Inc. youquan.li@stu.pku.edu.cn, zhouzenan@baichuan-inc.com, wentao.zhang@pku.edu.cn
1 Introduction
Equipped with advanced intelligence and formidable processing capabilities, large language models (LLMs) have demonstrated extensive potential in seamless interaction with human users and in assimilating real-time human feedback during inference processes (Fernandes et al., 2023). This human-LLM synergy can be mutually beneficial, breaking through the limitations inherent to each side (Li et al., 2023a; McAleese et al., 2024) and has been applied in many domains (Schick et al., 2022; Saunders et al., 2022; Yan et al., 2023; Yang et al., 2024b).
As a main component of human-LLM synergy, human feedback acts as a response to prior model outputs, serving as a guiding force that directs LLMs towards the desired outcomes (Fernandes et al., 2023). In practical applications, LLMs often need to iteratively adjust their responses based on user feedback in multi-turn dialogues to fulfill user needs. Effective feedback can enhance the quality of responses, while ineffective feedback may mislead LLMs. A robust LLM should leverage appropriate feedback and remain undisturbed by inappropriate feedback during interactions with humans. However, evaluating the responsiveness of LLMs to human feedback within multi-turn dialogues presents a significant challenge, as these models exhibit divergent behaviors compared to single-turn dialogues. As illustrated in Figure 1, LLMs that perform well in single-turn interactions may struggle to incorporate user feedback effectively. Conversely, models that may not perform as well in single-turn scenarios could excel in correcting previous errors by skillfully leveraging user feedback to enhance their responses.
Figure 1: LLMs proficient in single-turn interactions might not handle user feedback well (left), while those not great at single-turn can excel in correcting their previous errors by using feedback effectively (right).
Despite the advancement in evaluating LLMs, there remains a lack of a systematic benchmark for evaluating their responsiveness to various human feedback under real-world usage scenarios. Most existing benchmarks (Li et al., 2024; Liu et al., 2023a; Li et al., 2023b; Ni et al., 2024) assess LLMs within single-turn scenarios, which are naturally incapable of evaluating LLMs’ responsiveness to human feedback. Although there are benchmarks designed for multi-turn dialogues (Zheng et al., 2024; Sun et al., 2024; Bai et al., 2024), the user feedback is typically independent of the previous LLM output in a dialogue. While several benchmarks incorporating feedback exist (Wang et al., 2023; Yang et al., 2024a; Liu et al., 2023b), they typically assess LLMs on a singular task or within a specific domain, and the feedback is often not generated by real humans, failing to capture the nuance and diversity of human-LLM interaction.
In this work, we introduce FB-Bench, a fine-grained multi-task benchmark designed to evaluate LLMs’ responsiveness to various human feedback under real-world usage scenarios in Chinese. Drawing from the interaction scenarios of error correction and response maintenance, FB-Bench organizes a three-tier hierarchical taxonomy that encapsulates the fundamental elements of human-LLM interaction: user queries, model responses, and user feedback, as illustrated in Figure 2. It includes eight popular task types, five types of deficiencies in previous model outputs, and nine types of user feedback.
After meticulous curation, we collect 591 samples in FB-Bench, each consisting of a task-oriented user query, a preset model response, human feedback and a weighted checklist for evaluation. To precisely assess the performance of LLMs in a detailed manner, we employ GPT- 4o to act as a judge, scoring the model-generated follow-up responses based on the human-curated weighted checklist. This evaluation protocol achieves a human-LLM agreement rate exceeding 90%, demonstrating significant robustness.
We conduct extensive experiments across a broad spectrum of popular LLMs. The results indicate that the performance gap between open-source and closed-source LLMs is narrowing. Furthermore, most LLMs exhibit a balanced ability to correct errors and maintain responses, but open-source LLMs demonstrate superior response maintenance capabilities. We further analyze the impact of tasks, feedback, and previous responses’ deficiencies on LLMs’ responsiveness. These analyses reveals that
  • Leading LLMs show similar performance on each task in error correction. In contrast, their performance varies significantly in response maintenance.
  • Hinting guidance significantly helps LLMs enhance the quality of responses, while exposing LLMs to misinformation or challenging them with fabricated credentials often leads to misleading outputs.
  • Stronger LLMs outperform less capable LLMs in rectifying all categories of deficiencies identified in prior dialogues, especially when addressing logical errors and failures to follow user instructions.
To summarize, our work makes the following contributions:
  • New perspective. We develop a three-tier hierarchical taxonomy that encapsulates the fundamental elements of human-LLM interactions, focusing primarily on two main interactive scenarios: error correction and response maintenance.
  • New benchmark. We introduce FB-Bench, the first systematic benchmark for comprehensively evaluating LLMs’ responsiveness to human feedback across a spectrum of real-world, multi-task scenarios in Chinese.
  • More fine-grained evaluation. We develop a framework that employs a sample-specific weighted checklist to facilitate a fine-grained evaluation of each sample.
  • New findings. We perform a comprehensive evaluation of 27 different LLMs using FB-Bench, uncovering a significant performance discrepancy between error correction and response maintenance scenarios. We further analyze the factors that may impact the responsiveness of LLMs and provide valuable insights and directions for future research.
2 FB-Bench
In this section, we first outline the design logic behind FB-Bench in § 2.1 and  § 2.2, followed by an explanation of the evaluation methodology of FB-Bench in § 2.3. Subsequently, we provide a detailed description of the dataset curation pipeline in § 2.4 and finally present a statistical analysis of the dataset in §2.5.
Figure 2: Overview of FB-Bench. (1)Data Curation: A human-LLM synergy pipeline for mining target data from real-world scenarios and improving their quality and diversity. (2)Three-tier Hierarchical Taxonomy: Comprising 8 popular task types, 5 deficiency types and 9 feedback types, derived from two interaction scenarios. (3)Auto-Evaluation: A LLM-as-a-Judge framework to automatically evaluate LLM’s response with a weighted checklist.
2.1 Interaction scenario
In practical applications, error correction and response maintenance are two prevalent and significant scenarios. These scenarios encapsulate the essential dynamics between users and models, underscoring the importance of models’ ability to adapt and respond effectively to user feedback.
Error Correction: Users may pose a query and find the model’s response either objectively incorrect or unsatisfactory. Consequently, they provide feedback, expecting the model to acknowledge its response’s inadequacies and offer an improved version.
Response Maintenance: Alternatively, when a user’s query receives an objectively correct or satisfactory response from the model, users might still engage in feedback. This could be to either reaffirm or challenge the provided answer, aiming to verify the correctness and reliability of the information. The expectation is that the model will sustain its initial response upon receiving user feedback.
2.2 Hierarchical data taxonomy
A typical human-LLM interaction process comprises three components: the user’s query, the model’s response, and the user’s feedback. To ensure comprehensive coverage of various potential interaction scenarios and interaction types, we develop an extensive three-tier hierarchical taxonomy from the perspective of these three components.
2.2.1 Query task
From the perspective of user queries, the diversity of interactions primarily stems from the task type associated with each query. Therefore, we select eight popular tasks to encompass most real-world usage scenarios. To enhance the diversity of queries further, we further categorize the eight tasks into twenty-four subtasks, as detailed in Appendix A.1.1.
Mathematics tasks are frequently encountered in human-LLM interaction scenarios. Given the complexity of these problems, models often fail to provide accurate answers on their first attempt, necessitating collaboration between humans and models to resolve complex issues.
Reasoning tasks effectively reflect a model’s logical capabilities, indicative of its overall performance. Strong logical abilities enable the model to excel in other complex tasks, making it a vital component of human-LLM interaction.
Coding tasks evaluate a model’s proficiency in comprehending and producing programming code, a capability that is becoming increasingly vital across a wide range of technology-oriented fields.
Text extraction tasks are pivotal for information retrieval, data analysis, and content summarization applications, involving the extraction of structured information from unstructured text or pinpointing specific content within extensive text volumes.
Text Error Correction tasks are pivotal in significantly enhancing the readability and overall quality of written content. By fixing errors from typos to grammar, these tasks make text accurate and clear, highlighting their key role in keeping written communication professional and intact.
Text creation tasks not only test the model’s creativity and understanding but also play a crucial role in aiding people to express ideas more effectively and innovatively, enriching communication across various fields.
Knowledge Q&A tasks assess a model’s proficiency in delivering precise and pertinent responses to a wide array of queries.
Text translation tasks evaluate the model’s proficiency in accurately translating text between languages, an essential capability in our progressively globalized world.
2.2.2 Model response
From the perspective of the model’s response, it is either objectively correct or satisfies the user in response maintenance scenarios. In error correction scenarios, to enable more fine-grained research, we further categorize the deficiencies of model responses into the following five types:
  • Not following instructions: The response does not grasp or adhere to the given context, instructions, or format requirements.
  • Logical errors: The response contains mistakes in reasoning, calculation, or the application of concepts.
  • Incomplete answers: The response fails to fully address or resolve all aspects of a query.
  • Factual errors: The response includes incorrect or outdated information, encompassing grammatical and technical inaccuracies.
  • Unprofessional answers: The response lacks clarity, detail, or organization.
Figure 3: FB-Bench Statistics.
2.2.3 User feedback
From the perspective of user feedback, the interaction between humans and LLMs can be significantly influenced by the nature of the user feedback provided. We design a total of nine distinct types of feedback, comprising six for error correction and four for response maintenance, with one type overlapping between error correction and response maintenance. Table 1 provides a brief one-sentence description for each feedback within error correction and response maintenance scenarios.
Feedback Scenario Description
Pointing Out Errors EC Highlight specific inaccuracies or absurdities in the model’s output
Clarifying Intent EC Refine queries to guide the model towards more accurate and relevant responses.
Raising Objections EC Encourage the exploration of superior alternative solutions.
Detailed Explanation EC Request further information or a deeper understanding of the model’s response.
Hinting Guidance EC Guide the model at key points in problem-solving.
Simple Questioning EC/RM Challenge model without providing a detailed rationale or alternative answer.
Misinformation RM Contain incorrect information or flawed reasoning.
Credibility Support RM Challenge model’s response with fabricated authority or expertise.
Unreasonable Requests RM Propose demands or queries that fall outside ethical or common-sense boundaries.
Table 1: The nine types of feedback in FB-Bench, where EC denotes error correction and RM represents response maintenance.
2.3 Evaluation Protocol
Inspired by DRFR (Qin et al., 2024), we evaluate the quality of models’ follow-up responses by decomposing the evaluation criteria into a series of criteria that constitute a checklist. Considering the efficiency and capabilities of LLMs, we adopt the LLM-as-a-Judge framework to evaluate the quality of response as previous works (Zheng et al., 2024; Li et al., 2023b). Specifically, we employ GPT-4o to act as a judge, scoring the model-generated follow-up responses based on the human-curated checklist.
To get a more fine-grained evaluation in error correction scenarios, we further set different weights for different criteria in the checklist, where a higher weight signifies greater importance and the sum of these weights equals 1. If the response meets any criterion in the checklist, it receives the corresponding points. For $i$-th sample in error correction scenarios,
$score_{i}=\sum\limits_{j=1}^{n}w_{i,j}r_{i,j}$
where $w_{i,j}$ is the weight of $j$-th criterion, $r_{i,j}\in[0,1]$ denotes whether the $j$-th criterion within $i$-th sample is met.
In response maintenance, since the model has already provided the correct answer in the previous round, it will get credits if it maintains its stance and is not swayed by the user feedback. That’s to say, meeting any criterion in the checklist yields a score of 1.
$score_{i}=\begin{cases}1,&\forall{r_{i,j}=1},j\in[1,n]\\ 0,&\text{otherwise}.\end{cases}$
2.4 Dataset Curation
Each sample in FB-Bench mainly contains a task-oriented user query, a preset model response, human feedback and a human-curated weighted checklist for evaluation. The example can be found in Appendix A.1.3. The detailed construction pipeline is described as follows.
Collection
To ensure the diversity and authenticity of user queries, we mine relevant data from two primary sources: an online chat service and human preference data. Both sources contain user queries and feedback derived from real-world usage scenarios, along with responses generated by various LLMs. For error correction data, we employ heuristic rules to identify target data within the online chat service and select the response with the lowest score from human preference data. For response maintenance data, we adopt an opposite strategy to filter the target data from the two data sources. The detailed heuristic rules can be found in Appendix  A.2.1. After gathering the above data, we perform deduplication and anonymization, and categorize them into predefined tasks and subtasks using an in-house model to construct high task diversity data.
Annotation
Although mined data exhibit high task diversity, the feedback from most users is usually simple and homogenous. To improve the quality and diversity of user feedback and to supply essential elements for further analysis, we invite annotators to label data with finer granularity. Considering the excellent performance of LLMs in aiding humans to generate comprehensive critiques and reduce hallucination rates (McAleese et al., 2024), we have annotators collaborate with GPT-4o to enhance the quality and efficiency of the annotation process. Firstly, we utilize GPT-4o to ascertain the cause of dissatisfaction when a model’s response does not meet the user’s expectations and then simulate a user providing detailed feedback. Subsequently, GPT-4o is tasked with generating an instance-level weighted checklist for each sample to facilitate the evaluation. Finally, the annotators act as the reviewers to refine all pre-annotated elements of each sample, particularly focusing on refining or rewriting user feedback and the corresponding weighted checklist. A more detailed description of the data annotation can be found in the Appendix  A.2.2.
Figure 4: The subset evaluation results in FB-Bench between error correction and response maintenance scenarios. Overall denotes the mean of error correction score and response maintenance score. The dashed line represents the diagonal $y=x$.
Post-Filtering
To enhance distinguishment in scores among LLMs, we utilize three models, including Meta-Llama-3.1-8B-Instruct (Dubey et al., 2024), Phi-3-medium-4k-instruct (Abdin et al., 2024b), and Yi-1.5-9B-Chat (Young et al., 2024) as difficulty filters in our dataset curation pipeline. These models, with their diverse architectures and capabilities, provide a comprehensive assessment of the dataset’s difficulty. Specifically, we benchmark these models using this dataset, analyze their responses, and score them by GPT-4o. Finally, we discard samples for which all three models achieved full scores. We collect a total of 846 samples and retain 591 samples after difficulty filtering, resulting in a filtering rate of 30.14%.
2.5 Dataset Statistics
After meticulous curation, we collect 591 high-quality, diverse, and complex samples. The distributions of tasks, deficiencies in previous model responses within error correction scenarios, and user feedback within both error correction and response maintenance scenarios are all shown in Figure 3. More detailed statistics can be found in Appendix A.1.2
3 Experiments
3.1 Experimental Setup
Models. Given the considerations of performance, size and the popularity of LLMs, we systematically evaluate a wide array of LLMs, including GPT family, Claude-3.5, Qwen-2.5 family, ERNIE-4, Moonshot, Yi, Gemma-2, Mistral, InternLM, DeepSeek, GLM-4, Phi and LlaMa-3.1 family (Achiam et al., 2023; Team, 2024; moo, ; Team et al., 2024; Jiang et al., 2023; Cai et al., 2024; Liu et al., 2024; GLM et al., 2024; Abdin et al., 2024b, a; Dubey et al., 2024).
Response generation. We employ the official settings and chat template in HuggingFace model card for open-source LLMs. Proprietary models are assessed via their official API endpoints. Considering the varied requirements for diversity and creativity across tasks, we set different temperatures for different tasks. More details can be found in Appendix A.3.1.
Evaluation. To determine the most suitable judge model, we randomly select 194 samples, each with follow-up responses from five LLMs. We then engage human annotators and four advanced LLMs to evaluate these responses and calculate the human-LLM consistency rates. We found that gpt-4o-2024-08-06 achieved the highest consistency rate, at 90.91%. The detailed experimental setup and results can be found in Appendix A.4. Consequently, we choose gpt-4o-2024-08-06 as judge model to evaluate each generated follow-up response based on the corresponding weighted checklist. To enhance the determinism of the judgment, we set the temperature to 0 and the output length to 4096. The evaluation prompt and cases are presented in Appendix A.3.2.
3.2 Main Results
The subset evaluation results in FB-Bench are presented in Figure 4, with detailed results available in Appendix A.4.1. The main findings are as follows:
The performance gap between open-source LLMs and closed-source LLMs is narrowing. DeepSeek-V3 ranks first in overall performance, followed by qwen-max-0919 and gpt-4o-2024-11-20, demonstrating the best error correction ability and third in response maintenance ability. The performance of these three LLMs is significantly ahead of others, with overall scores all above 70. The fourth-ranked LLM, Qwen2.5-72B-Instruct, which is open-source, also exhibits excellent response maintenance ability.
Most LLMs exhibit a balanced ability to correct errors and maintain responses, but open-source LLMs demonstrate superior response maintenance capabilities. Most models are positioned near the diagonal $y=x$, indicating their balanced proficiency in error correction and response maintenance. However, some open-source LLMs are situated above the diagonal $y=x$, suggesting they excel more in maintaining responses than in correcting errors. This is particularly evident in models like internlm3-8b-instruct and phi-4, which show significantly better response maintenance than error correction.
3.3 Analysis
Thanks to our comprehensive taxonomy, we can delve into several critical factors that significantly influence the performance of LLMs on FB-Bench, including task types, feedback types and deficiency types.
Leading LLMs show similar performance on each task in error correction. In contrast, their performance varies significantly in response maintenance.
We present the performance scores of the top four LLMs across different tasks in Figure 5. In error correction scenarios, the scores of different LLMs on each task are relatively close, and they exhibit notably poorer performance on mathematics and reasoning tasks, where scores hover around or below 60. Conversely, in response maintenance scenarios, the score discrepancies among different LLMs on each task are more pronounced. Specifically, in reasoning tasks, the performances of the four LLMs vary significantly, with gpt-4o-2024-11-20 lagging considerably behind the other three. Additionally, Qwen-2.5-72B-Instruct also falls significantly behind the other three LLMs in text translation tasks.
Figure 5: The performance of top four LLMs across eight popular tasks
Hinting guidance significantly helps LLMs enhance the quality of responses, while exposing LLMs to misinformation or challenging them with fabricated credentials often leads to misleading outputs.
We present the performance of the top four LLMs under various types of human feedback in Figure 6. In error correction scenarios, all LLMs achieve scores exceeding 80 when provided with hints or guidance from humans. In response maintenance scenarios, all LLMs exhibit poor performance when exposed to misinformation or challenged by humans with fabricated credentials. Furthermore, it is evident that qwen-max-0919 significantly outperforms all other LLMs when handling unreasonable requests, indicating its superior safety capabilities.
Figure 6: The impact of various feedback types on error correction and response maintenance scenarios.
Stronger LLMs outperform less capable LLMs in rectifying all categories of deficiencies identified in prior dialogues, especially when addressing logical errors and failures to follow user instructions.
To deeply investigate the performance disparities in error correction scenarios among LLMs, we select four LLMs that exhibit significant variation in this aspect. Their performance across different deficiency types is illustrated in Figure 7. The results indicate that stronger LLMs consistently outperform less capable ones across all deficiency categories. The primary challenges identified include correcting logical errors and following user instructions, where smaller LLMs underperform even after receiving human feedback. This underperformance is likely attributable to the limited capabilities of smaller LLMs.
Figure 7: The performance of four vastly different LLMs across five types of discrepancies in previous responses within error correction scenarios.
4 Related Work
Evaluation of LLMs
The evaluation of LLMs is essential for their development. It reveals the strengths and weaknesses of existing models and offers key insights and directions for future research. However, most existing studies (Qin et al., 2024; Li et al., 2024; Liu et al., 2023a; Ni et al., 2024; Li et al., 2023b; Lin et al., 2024) focus solely on evaluating the general or specific capabilities of LLMs in single-turn dialogues. They fail to assess LLM performance under various user feedback, which typically involves multi-turn dialogue scenarios. Although there are some benchmarks for multi-turn LLMs (Zheng et al., 2024; Sun et al., 2024; Bai et al., 2024; Kwan et al., 2024; Liang et al., 2024), the user inputs in the multi-turn dialogues are often independent, lacking feedback towards to the previous LLM output. Furthermore, much of the data in these multi-turn dialogue benchmarks is synthesized by LLMs, failing to exhibit the diversity and complexity of real-world scenarios.
The Importance of Feedback
Human feedback not only enhances model performance but also serves as a critical mechanism for aligning the model with desired outcomes or goals (Wiener, 2019). Training models on feedback data, not only can directly enhance the quality of the generated content (Ouyang et al., 2022) but also allows models to better align with human preferences in style and tone (Ziegler et al., 2019). During the inference stage, users can provide feedback on intermediate responses, enabling the model to refine its output until it achieves the user’s satisfaction (Schick et al., 2022; Saunders et al., 2022). However, a systematic benchmark for evaluating the impact of human feedback on LLMs during the inference stage is still lacking.
Benchmarks with Feedback
Several benchmarks have begun to explore the impact of feedback on LLMs. However, they predominantly focus on specific tasks or domains. MINT (Wang et al., 2023) exclusively assesses the coding and reasoning capabilities of LLMs that utilize tools and receive AI-generated language feedback. Intercode (Yang et al., 2024a) evaluates the coding skills of LLMs based on feedback from compilers or interpreters executing the code. AgentBench (Liu et al., 2023b) examines the reasoning and decision-making abilities of LLMs-as-Agents in response to environmental feedback. Different from prior works, FB-Bench introduces a novel approach by measuring the responsiveness of LLMs to diverse user feedback across a broad spectrum of real-world usage scenarios.
5 Conclusion
We introduce FB-Bench, a fine-grained muti-task benchmark for comprehensively evaluating the responsiveness of LLMs to various human feedback across real-world usage scenarios in Chinese. A three-tier hierarchical taxonomy, grounded in two human-LLM interaction scenarios, is established to ensure thorough coverage of diverse interaction types and scenarios. To facilitate a fine-grained and accurate evaluation, a LLM-as-a-Judge framework, equipped with a weighted checklist, is employed. Benchmarking results from 27 well-known LLMs demonstrate significant performance variations between error correction and response maintenance. Further analysis explores the principal factors influencing the responsiveness of LLMs and provides valuable insights for subsequent research.
Limitations
Here we discuss several limitations of this work.
  • Although our data curation pipeline and evaluation framework are designed to be flexible and adaptable for use across various languages, the FB-Bench dataset is currently available only in Chinese. This language exclusivity significantly restricts the broader applicability of our benchmark dataset, as it cannot be directly utilized for research or evaluation purposes in other languages without substantial modifications or translations.
  • Our evaluation protocol primarily relies on LLMs. Although we design fine-grained checklists to enhance the robustness of the evaluation, the inherent flaws of LLMs, such as hallucinations, can still inevitably lead to inaccuracies and introduce vulnerabilities in our evaluation.
Appendix A Appendix
A.1 Detailed description of the dataset
A.1.1 Query subtask
We further categorize the eight tasks into twenty-four subtasks. Table 2 presents a brief one-sentence description of each subtask.
Task Subtask Description
Mathematics Algebra Solving algebraic expressions and equations.
Geometry Understanding and applying geometric principles and theorems.
Equations and inequalities Solving for variables within equations and inequalities.
Combinatorial probability Calculating the likelihood of various combinations and outcomes.
Arithmetic Performing basic mathematical operations and number theory.
Reasoning Common sense reasoning Applying everyday knowledge and logic to solve problems.
IQ questions Solving puzzles and questions designed to measure intelligence.
Coding Code generation Automatically writing code snippets for given tasks.
Code debugging Identifying and fixing errors or bugs in code.
Code knowledge Understanding programming concepts, languages, and frameworks.
Text extraction Information extraction Extracting structured information from unstructured text.
Summary generation Creating concise summaries of lengthy texts.
Title extraction Identifying and extracting the principal titles or headings from documents
Text error correction Typo detection Identifying and correcting misspelled words in the provided text.
Text proofreading Examining texts for errors in logic, factuality, or coherence
Grammar checking Identifying and rectifying grammatical errors
Text creation Style-based rewriting Adapting content to different tones, styles, or formats.
Generation Producing coherent, contextually relevant content from scratch.
Knowledge Q&A Objective facts Q&A Providing answers to questions based on factual information.
Conceptual explanation Explaining theories, concepts, or ideas in a comprehensible manner.
Experiential advice Offering advice based on personal or shared experiences.
Logical reasoning Applying logic to solve problems or answer questions.
Text translation Chinese to English Translating text from Chinese to English accurately.
English to Chinese Translating text from English to Chinese accurately.
Table 2: The description of twenty-four subtasks across eight tasks in FB-Bench.
A.1.2 Detailed data statistics
The distribution of task and subtask categories in our dataset is shown in Figure 8.
Figure 8: The distribution of task and subtask categories in FB-Bench.
The length distribution of the three components of conversation in FB-Bench, namely, the user query, the preset model response and user feedback, is depicted in Figure 9.
Figure 9: The length distribution of three components in conversations within FB-Bench.
A.1.3 Example in FB-Bench
We select examples from the error correction and response maintenance scenarios, which are displayed in Figure 10.
Figure 10: Examples wthin error correction and response maintenance scenarios in FB-Bench.
A.2 Detailed description of data curation
A.2.1 Heuristic rules
For error correction data, we initially filter target data from online chat data using the following heuristic rules:
  • Multi-turn dialogues
  • Created within the past three months
  • The model’s second-turn response contains phrases such as “I’m sorry”, “I apologize”, or “Please forgive me”.
  • Token length is less than 5000
  • Does not use retrieval augmentation
  • No security restrictions are triggered
Subsequently, we include samples with scores below 3 from human preference data, where the model’s responses are deemed poor in quality. Scores for human preference data range from 1 to 5, with higher scores indicating better quality.
For response maintenance data, we include answers and corresponding dialogues that have received user upvotes by analyzing the front-end log records. Additionally, we remain data with scores above 4 from the human preference dataset.
A.2.2 Data annotation
We employ a two-stage labeling process, which includes GPT pre-annotation followed by human annotation, to label each sample in the collected dataset.
GPT pre-annotation
For each sample within error correction scenario, GPT-4o-2024-05-13(denoted as GPT-4o for convenience in the following text) is utilized to generate a series of preliminary labels. This process involves using GPT-4o to produce a reference response to the user’s query. Subsequently, GPT-4o identifies the reason for user dissatisfaction and simulates a user providing detailed feedback, assigning a corresponding predefined feedback category. Following this, GPT-4o generates a follow-up response based on the user query, the model’s response, and the feedback provided. Finally, GPT-4o generates an instance-level checklist consisting of several yes/no questions, each associated with a weight ranging from 0 to 1, where a higher weight signifies greater importance. The cumulative weight of these questions equals 1. For each sample within response maintenance scenario, since the model response is good enough, we employ GPT-4o to simulate a real user providing several detailed feedback for each of the predefined feedback types. We then ask GPT-4o to generate an instance-level checklist, which also consists of several yes/no questions, albeit without associated weights.
Human annotation
To further improve the diversity and quality of the dataset, we engage ten Chinese annotators, each holding a bachelor’s degree, to review, revise, and refine the data initially annotated by GPT-4o. To provide guidance on data annotation, we initially prepare a document detailing the dataset’s background, the meaning of each field, the value ranges, and specific examples. Subsequently, we organize the data into a table and annotate three entries for each task. Finally, we conduct a conference to instruct human annotators on data annotation, integrating the document and sample data. Following the guidance, human annotators are asked to randomly select and annotate five entries per task. We then review these annotations to ensure quality and consistency, confirming their understanding of our requirements. In the labeling process for each sample in the error correction scenario, human annotators first assess whether the task, subtask, and deficiency types are misaligned. They then refine or rewrite the human feedback and its corresponding feedback type, followed by refining or rewriting the associated checklist. This process involves modifying all content related to yes/no questions and the corresponding weights. For each sample in the response maintenance scenario, annotators similarly begin by checking the task and subtask types. They then provide the human feedback and its corresponding feedback type, concluding with the checklist, as in the error correction scenario. Each sample in FB-Bench is annotated by one annotator and subsequently reviewed by another.
A.3 Detailed description of experiments
A.3.1 Response generation
We utilize the vllm library (Kwon et al., 2023) to deploy open-source LLMs for generating follow-up responses based on a user query, a preset model response, and human feedback. In terms of temperature settings, we assign distinct values for different tasks: 0.7 for text creation and text translation, 0.1 for knowledge-based Q&A, and 0 for all other tasks. For the maximum output length, we set it to the minimum value between 4096 and the difference between the LLM context length and the context tokens length.
A.3.2 Evaluation
Since FB-Bench focuses on Chinese, we employ a Chinese prompt for evaluation, which is present in Figure 11.
Figure 11: The evaluation prompt template of FB-Bench.
Figure 12 and Figure 13 show the corresponding evaluation cases within error correction and response maintenance scenarios, respectively.
Figure 12: The evaluation case in error correction scenario.
Figure 13: The evaluation case in response maintenance scenario.
A.4 Agreement Evaluation
To justify the weighted checklist LLM-as-a-Judge method we designed for FB-Bench, we conduct extensive human evaluation over FB-Bench’s randomly selected examples.
We collect 194 samples by performing random sampling for each task within FB-Bench, along with the following-up responses from five LLMs: claude-3-5-sonnet-20240620, gpt-4o-2024-05-13, qwen-max-0919, Qwen2.5-72B-Instruct, and internlm2_5-20b-chat. We then replace the selected models’ names with Model1 to Model5 and ask human annotators to evaluate the following-up responses based on the corresponding checklist. Subsequently, we employ several advanced LLMs, including including gpt-4o-2024-08-06, gpt-4o-2024-11-20, qwen-max-0919, and claude-3-5-sonnet-20241022, as judge models to perform evaluations as well. Finally, we compile all the judgment results generated by human annotators and the judge models to calculate the consistency rate.
Table 4 presents the consistency rate between human annotators and several judge models. The consistency rate for gpt-4o-2024-08-06 is the highest, at 84.79%. Consequently, we select gpt-4o-2024-08-06 as the judge model in FB-Bench. The corresponding score of the five selected LLMs under different judge models can be found in Table 5.
A.4.1 The full results in FB-Bench
We evaluate 27 popular LLMs using FB-Bench, with the results presented in Table 3.
Model Error Correction Response Maintenance Overall
DeepSeek-V3 74.24 75.48 74.86
qwen-max-0919 72.03 74.42 73.22
gpt-4o-2024-11-20 71.74 73.68 72.71
Qwen2.5-72B-Instruct 67.95 71.29 69.62
ERNIE-4.0-8K-0329 62.37 75.85 69.11
claude-3-5-sonnet-20241022 69.92 67.31 68.61
Mistral-Large-Instruct-2411 64.97 70.36 67.66
yi-lightning 70.09 64.69 67.39
glm-4-0520 60.25 73.43 66.84
gpt-4o-2024-05-13 68.82 64.72 66.77
claude-3-5-sonnet-20240620 68.99 63.95 66.47
phi-4 58.17 74.67 66.42
internlm3-8b-instruct 50.85 81.04 65.95
moonshot-v1-32k 59.09 69.27 64.18
gpt-4o-mini-2024-07-18 61.83 63.74 62.79
internlm2_5-20b-chat 53.44 71.79 62.61
Qwen2.5-7B-Instruct 57.09 58.62 57.85
internlm2_5-7b-chat 48.14 67.45 57.79
glm-4-9b-chat 58.60 54.92 56.76
Yi-1.5-34B-Chat 46.98 60.26 53.62
Meta-Llama-3.1-70B-Instruct 54.51 52.69 53.60
Phi-3-medium-4k-instruct 37.62 65.54 51.58
Ministral-8B-Instruct-2410 49.77 51.55 50.66
gemma-2-27b-it 48.55 51.24 49.89
Yi-1.5-9B-Chat 43.62 54.39 49.00
gemma-2-9b-it 44.56 46.84 45.70
Meta-Llama-3.1-8B-Instruct 51.17 38.21 44.69
Table 3: The full evaluation results in FB-Bench between error correction and response maintenance scenarios. The bold, underlined, and tilde denote the first, second, and third rankings, respectively.
Judge Model Consistency Rate
gpt-4o-2024-08-06 90.91%
gpt-4o-2024-11-20 88.77%
qwen-max-0919 89.42%
claude-3-5-sonnet-20241022 90.47%
Table 4: The consistency rate between human and different judge models.
gpt-4o-2024-05-13 claude-3-5-sonnet-20240620 internlm2_5-20b-chat Qwen2.5-72B-Instruct qwen-max-0919
gpt-4o-2024-08-06 75.11 73.88 72.28 70.71 66.59
gpt-4o-2024-11-20 74.61 75.79 71.99 69.99 60.62
qwen-max-0919 77.15 77.76 73.04 75.05 66.36
claude-3-5-sonnet-20241022 73.96 76.07 71.02 72.59 62.42
Table 5: The overall score of the five selected LLMs under different judge models. The header represents the evaluated models, while the index indicates the judges.